### v1

@@ Abstract
Generalizable manipulation of articulated objects remains a challenging problem in many real-world scenarios given the diverse object structures, functionalities, and goals. In these tasks, both semantic interpretations and physical plausibilities are crucial for a policy to succeed. To address this problem, we propose SAGE, a novel framework that bridges the understanding of semantic and actionable parts of articulated objects to achieve generalizable manipulation under language instructions. Given a manipulation goal specified by natural language, an instruction interpreter with Large Language Models (LLMs) first translates them into programmatic actions on the object’s semantic parts. This process also involves a scene context parser for understanding the visual inputs, which is designed to generate scene descriptions with both rich information and accurate interaction-related facts by joining the forces of generalist Visual-Language Models (VLMs) and domain-specialist part perception models. To further convert the action programs into executable policies, a part grounding module then maps the object semantic parts suggested by the instruction interpreter into so-called Generalizable Actionable Parts (GAParts). Finally, an interactive feedback module is incorporated to respond to failures, which greatly increases the robustness of the overall framework. Experiments both in simulation environments and on real robots show that our framework can handle a large variety of articulated objects with diverse language-instructed goals. We also provide a new benchmark for language-guided articulated-object manipulation in realistic scenarios. More details and demonstrations can be found at https://geometry.stanford.edu/projects/sage/.
@@ I Introduction
Generalizable object manipulation, which involves the transfer of manipulation skills across object shapes [38, 21], object categories[14, 13, 57, 53], or even tasks [61], is gaining increasing attention in the robot learning community. A line of works has focused on learning generalizable atomic manipulation skills of articulated objects [38, 21, 14, 13, 17, 15]. Yet this remains a challenging problem, especially in real-world scenarios with complex object structures (such as home appliances) and long-horizon tasks specified by natural language instructions.
Imagine that you bought a new blender for your kitchen. You may play it with its power off, pressing its buttons, opening its lid, or taking off its container. But to make juice with it, you need to truly understand the functionalities of these parts, for example, “pressing the container turns it on, and it will mince the contents in the container”. In Gibson’s theory of distributed cognition [62], your interactions with the blender without and with prior knowledge can be explained by two different types of affordances: physical affordance and cognitive affordance. Physical affordances are for tasks that are mainly constrained by physical structures, such as the part and joint structures of articulated objects. Cognitive affordances, on the other hand, are provided by cultural conventions and involve a high-level semantic understanding of the object.
@@
To smartly interact with varying types of articulated objects, a general framework is needed for generating and executing actions that are both semantically meaningful and physically plausible. To this end, we introduce SAGE, a framework that bridges the semantic and actionable part understandings of articulated objects to achieve generalizable manipulations under natural language instructions (Fig. LABEL:fig:teaser). Given the RGBD scan of an articulated object and a language-specified manipulation goal as input, we design an instruction interpreter with Large Language Models (LLMs) to translate the language instruction into action programs comprised of so-called action units, which are represented by 3-tuples of semantic parts, joint types, and state changes. This process also involves an image caption module applied to the RGB inputs to convert the visual contexts into language descriptions that can be understood by the LLMs. Combining large generalist Visual-Language Models (VLMs) and domain-specialist part perception models, we are able to acquire scene descriptions that are both informative and accurate about interaction-related facts. We then convert the action programs defined on the object semantic parts into executable policies through a part grounding module that maps the semantic parts into so-called Generalizable Actionable Parts (GAParts), which is a cross-category definition of parts according to their actionabilities. To achieve this, we employ a large open-vocabulary 2D segmentation model and again combine it with a domain-specialist GAPart model by merging their part proposals and ground them together to GAPart labels. From the detected GAParts, we are able to generate physically plausible part actions indicated by the pre-obtained action programs. Finally, we also introduce an interactive feedback module that actively responds to failed action steps and adjusts the overall policy accordingly, which enables the framework to act robustly under environmental ambiguities or its own failure.
@@
We demonstrate the effectiveness of our method both in simulation environments and on real robots. Specifically, we introduce three novel benchmarks to assess crucial components of language-guided articulated-object manipulation: scene descriptions, part perceptions, and manipulation policies. Quantitatively, our method surpasses all baseline approaches. This success is attributed to our blend of a comprehensive general-purpose model and a precise domain specialist, endowing our method with both stellar performance and robust generalization capabilities. Qualitatively, we further illustrate the proficiency of our method through its performance on several demanding tasks.
@@
  • We introduce a novel perspective of bridging semantic and actionable parts and construct a robot system for generalizable manipulation of articulated objects.
  • We explore the idea of joining the forces of general-purpose large models and domain-specific models extensively throughout our pipeline from scene understanding, part perception, to manipulation.
@@
  • We also provide a new benchmark for articulated-object manipulation including evaluations of scene descriptions, part perceptions, and manipulation policies.
@@ II Related Work
Articulated-object manipulation. Articulated object manipulation presents significant challenges due to diverse object geometries and physical characteristics. Although benchmarks by [51, 38] have been introduced, their scope remains limited. While methods like [41, 26, 4] have delved into motion planning, visual affordance learning has also gained attention [37, 55, 54, 63, 15]. Other works [7, 58] have proposed special representations for manipulation, though they are tailored mainly for suction grippers. On the other hand, [14, 13] introduce generalizable parts as a novel representation but are limited to single-part interaction and can not handle natural instructions.
@@
Generalizable Object Manipulation. Achieving generalization in robot applications is essential yet challenging. Various works [8, 49, 18, 12, 59] employ supervised learning with motion planning to tackle generalizable tasks such as grasping. Nevertheless, these techniques often necessitate task-specific architectures and may falter in intricate manipulations. Though reinforcement learning holds promise for complex challenges [42, 2], the lack of generalizability remains unresolved [31, 16]. The ManiSkill benchmark [38] pioneers category-level object manipulation in simulations, while [44] utilizes imitation learning with RL-trained demonstrations. However, in both cases, the generalization ability is still unclear. [14, 13] first tackles generalizable manipulation in a cross-category manner, but is still limited to predefined part classes. [57, 53] have universal generalization ability but only for grasping tasks. In summary, Generalizable object manipulation is still extremely challenging.
@@
Language models for robotics. Incorporating pre-trained language models into robot systems has been gaining tremendous attention recently. A notable segment of these works concentrates on the grounding of language to manipulation skills, essential for executing long-horizon instructions [27, 48, 36, 5, 43, 1, 17, 19, 45, 22, 60]. A hierarchical approach, seen in another subset of studies, hinges on a two-fold process: establishing a skill set and subsequently strategizing over it using large language models (LLMs) [3, 25, 47, 23, 33, 29]. However, this method is limited by the huge data collection effort, is still constrained to the predefined skill sets, and lacks visual understanding. In contrast, LLMs are employed to spawn code for low-level skills through APIs, though predicting these skills’ continuous parameters remains challenging [32]. Some employ a way-point-based method for efficiency, sidestepping low-level robot actions in favor of trajectory hand poses [46, 17]. But they also require lots of data and have no detailed planning strategy. VoxPoser [24] performs motion planning based on affordance and obstacle maps generated by LLM, but still lacks visual understanding. For articulated objects with diverse geometries, intricate structures, and physical constraints, language-guided manipulation remains a challenging and underexplored problem.
III Method
@@
Figure 2: Context-aware instruction interpretation. With instruction and observation (RGBD image) as input, the interpreter first generates a scene description with VLM and GAPartNet[14]. Then LLM(GPT-4) takes both instruction and scene description as input and generates semantic parts and action programs. Optionally, we can input a specific user manual, and LLM will generate a target actionable part; see Sec.III-B for more details about the actionable part.
@@
Figure 3: Scene description. To better help with action generation, the scene description should contain object information, part information, and some interaction-related information. We use expert GAPart model[10] to generate some expert descriptions as part of the VLM prompt and then generate the scene descriptions, which works well and absorb the advantages of both models
III-A Instruction interpretation
Action programs
@@ III-A Instruction interpretation
The goal of the instruction interpreter is to translate natural-language instruction $\tilde{I}$ into programmatic instructions. We define the most basic manipulation on one single part of articulated objects as an “action unit”, which can be represented by a 3-tuple ${\bm{a}}=(\tilde{p},j,\Delta s)$ with part name $\tilde{p}$, joint type $j$, and the change of part state $\Delta s$. The part name $\tilde{p}$ is a noun in natural language which will later be sent into an open-vocabulary segmentation module; $j$ refers to the joint directly linked to part $\tilde{p}$ and indicates whether it is a revolute or a prismatic joint; $\Delta s$ is the state change of joint $j$ under the action, which is either an angle ±$\pm\theta$ if $j$ is revolute or a relative translation ±$\pm t$ w.r.t. the part bounding box if $j$ is prismatic. The final output of the interpreter should be in one of the following formats:
@@
Here the Union expression is for non-deterministic policy generation when multiple action units can reach the same goal, e.g., both pulling the door and pressing a button result in the microwave door being opened. List is for sequential action generation when no single-step solution exists, e.g., in order to open a door, a knob must be first rotated to unlock the door.
Visual input parsing
We also incorporate a visual input parser into our instruction interpretation process, which translates the RGB inputs into language descriptions that can be understood by LLMs. A naive solution is to directly employ a large Visual-Language Model (VLM), such as Google Bard and GPT-4V, to generate scene captions. However, we notice that most of these models lack the capability to provide exact task-related facts. Inspired by [20], we recognize that the inclusion of expert facts can enhance the correctness of visual perception. Thus, we first detect the actionable parts using the GAPartNet [14] segmentation model for actionable parts and then pass it into the VLM together with the task description to generate a comprehensive scene description. Combining the expert knowledge of the specialist model and the general knowledge of VLM, we can acquire scene descriptions that are more accurate and also task-related (Fig. 3).
Language instruction interpretation
We employ GPT-4 [39], an LLM, to parse the input instruction, which outlines the intended goal. Here is an example of the overall inputs and outputs of the instruction interpreter: Certain objects, such as appliances, have very diverse and intricate structures that can hardly be covered by general knowledge. Therefore, we also accommodate the additional input of the object usage manuals, which can be easily accessed together with the purchase of the object in real-life applications. This feature allows the LLM to gain expert knowledge of the object and thereby provide more precise and relevant outputs.
III-B Part Grounding
@@
After getting the results from language instruction interpretation, which includes the semantic part information and corresponding semantic-level interaction strategy, we then first ground the semantic part to the physical part. In particular, with the semantic part, e.g. handle as input, we ground it to an actionable part with its pose.
Part discovery. In our methodology for part-level object perception, we leverage the strengths of both broad-spectrum 2D vision models and specialized 3D models.
Initially, we deploy the 3D GAPart Model [14] to enhance perception. This model interprets the partial RGBD point cloud $X$ to yield its set of 3D part proposals, which are masked pointclouds denoted as $\mathcal{M}_{j}^{3D}=\{M_{j}^{3D}\}$. Subsequently, our general-purpose large model processes the RGB image $I_{rgb}$ in tandem with the semantic parts $\{\tilde{p}\}$ from action unit ${\bm{a}}$, producing a series of 2D part proposals represented as $\mathcal{M}_{j}^{2D}=\{M_{j}^{2D}\}$. By synergizing the robust capabilities of general-purpose 2D vision models, such as SAM [30] and GroundedDINO [34], with the expertise of the domain-specific 3D GAPart Model, we achieve a comprehensive and nuanced part perception.
Actionable part grounding. Then we introduce an actionable part grounding model, which takes both the 2D and 3D part proposals as input and outputs the grounded GAPart labels. To this end, We first collect a dataset of actionable part features and part labels $\{{F}^{\text{dino}}_{\text{part}},l_{\text{part}}\}$. In particular, we extract a good image feature map $F^{\text{dino}}$ from DINOv2 [40]. Then, for the target actionable part denoted as $p_{j}$, we use its mask to get the part feature $F^{\text{dino}}_{j}$. We first do max-pooling to obtain the part feature and run KNN algorithm to find the grounded actionable part label denoted as $l_{j}$
@@
Actionable part estimation. We introduce a pose and joint estimation model to perceive grounded actionable parts. Similar to GAPartNet, the pose estimation model takes the point cloud $X$ as input and estimates the 6D part pose. Our model tackles part symmetry similar to GAPartNet, and we can directly infer the interaction joint state from the part pose thanks to the GAPart definition.
III-C Action Generation
Once we have grounded the semantic part to the actionable part, we can generate executable manipulations on this part. We first estimate the part pose as defined in GAPartNet [14]. We also compute the joint state (part axis and position) and plausible motion direction based on the joint type (prismatic or revolute). Then we generate the part actions according to these estimations. We first predict an initial gripper pose as the primary action. It is followed by a motion induced from a pre-determined strategy defined in GAPartNet [14] according to the part poses and joint states. For example, to open a door with a revolute joint, the starting position can be on the door’s rim or handle, with the trajectory being a circular arc oriented along the door’s hinge.
III-D Interactive Feedback
Algorithm 1 Interactive perception
      Output joint axis $j_{t}$ and movement $s_{t}$
2:$j_{t},s_{t}=\text{GeometryInference}(R_{t},t_{t})$
Algorithm 2 Tracking and planning update

### v2

@@ Abstract
To interact with daily-life articulated objects of diverse structures and functionalities, understanding the object parts plays a central role in both user instruction comprehension and task execution. However, the possible discordance between the semantic meaning and physics functionalities of the parts poses a challenge for designing a general system. To address this problem, we propose SAGE, a novel framework that bridges semantic and actionable parts of articulated objects to achieve generalizable manipulation under natural language instructions. More concretely, given an articulated object, we first observe all the semantic parts on it, conditioned on which an instruction interpreter proposes possible action programs that concretize the natural language instruction. Then, a part-grounding module maps the semantic parts into so-called Generalizable Actionable Parts (GAParts), which inherently carry information about part motion. End-effector trajectories are predicted on the GAParts, which, together with the action program, form an executable policy. Additionally, an interactive feedback module is incorporated to respond to failures, which closes the loop and increases the robustness of the overall framework. Key to the success of our framework is the joint proposal and knowledge fusion between a large vision-language model (VLM) and a small domain-specific model for both context comprehension and part perception, with the former providing general intuitions and the latter serving as expert facts. Both simulation and real-robot experiments show our effectiveness in handling a large variety of articulated objects with diverse language-instructed goals.
Figure 1: Overview. We present SAGE, a framework bridging the understanding of semantic and actionable parts for generalizable manipulation of articulated objects. Left: We give two examples of human instructions illustrating the concept of part semantics, part functionalities, and corresponding actions. Right: examples of real-world tasks and our results. Our framework can tackle diverse manipulation tasks on various articulated objects.
@@ I Introduction
From furniture to home appliances, articulated object prevails in our daily lives. While being long studied in the robotics literature, their diverse object structures, functionalities, and manipulation goals in real-world scenarios remain challenging for most robot systems. The design of a general system that accommodates such object diversities falls into the problem of generalizable object manipulation, which involves the transfer of manipulation skills across object shapes [36, 20], object categories[13, 14, 56, 51], or even tasks [60]. Specifically, a line of works has focused on learning generalizable manipulation skills of articulated objects [36, 20, 13, 14, 17, 15]. However, even these works are still limited to a few common object categories, leaving aside the problem of how to model the resemblance between seemingly irrelevant articulated objects, from a simple cabinet to a multifunctional blender with intricate mechanical structures and electricity-powered functionalities.
In this work, we approach this problem by separately modeling the cross-category part commonality in their semantics and actions. Taking the blender in Fig. 1 top left as an example, to mince the food ingredients in it, one needs to press its container from the top to turn it on. The part you press on is semantically a “container”, but considering your interaction with it, its action resembles a “button” which can be pressed and trigger other mechanical motions (such as the blade rotating). To exert the functionality of this object, both part semantics and actions should be well understood. This also fits into the two different types of affordances in Gibson’s theory of distributed cognition [61]: physical affordance and cognitive affordance, which have different emphases in human cognition of interactable objects. Physical affordance focuses on physical structures such as the joint conditions and part motions of articulated objects; cognitive affordances, on the other hand, are provided by cultural conventions and involve a high-level semantic understanding of the object.
Following this inspiration, we introduce SAGE, a framework for generalizable manipulation of articulated objects under natural language instructions by bridging semantic and actionable parts (Fig. 1 right). Our key insight is that large Visual-Language Models (VLMs) possess general knowledge of part semantics, while small domain-specific models present higher accuracy in predicting part actions, which can serve as “expert facts”. Different from prior works that separately assign VLMs and small models to different sub-tasks[24], we fuse their predictions in both context comprehension and part perception, which achieves a good balance of generality and exactness.
@@
Concretely, given a manipulation goal specified by natural language and an RGBD image as visual observation, an instruction interpreter first translates the language instruction into programmatic action representation. These action programs are composed of so-called action units, which are represented by 3-tuples of object semantic parts, joint types, and state changes. We then convert the action programs defined on the object semantic parts into executable policies through a part grounding module that maps the semantic parts into so-called Generalizable Actionable Parts (GAParts), which is a cross-category definition of parts according to their actionabilities. From the detected GAParts, we can generate physically plausible part actions indicated by the pre-obtained action programs. Finally, we also introduce an interactive feedback module that actively responds to failed action steps and adjusts the overall policy accordingly, which enables the framework to act robustly under environmental ambiguities or failure.
@@
We demonstrate the effectiveness of our method both in simulation environments and on real robots. Specifically, we also assess the crucial components of language-guided articulated-object manipulation: scene descriptions, part perceptions, and manipulation policies. Quantitatively, our method surpasses all baseline approaches. This success is attributed to our blend of a comprehensive general-purpose model and a precise domain specialist, endowing our method with both stellar performance and robust generalization capabilities. Qualitatively, we further illustrate the proficiency of our method through its performance on several demanding tasks.
@@
  • We bridge the notion of semantic and actionable parts to model the commonality across articulated objects of highly diverse structures and functionalities.
  • We build a robot system for generalizable manipulation of articulated objects under language instructions.
  • We design a knowledge fusion mechanism between VLMs and small domain-specific models to incorporate expert facts into general perception and comprehension.
@@ II Related Work
Articulated object manipulation presents significant challenges due to diverse object geometries and physical characteristics. Although benchmarks by [49, 36] have been introduced, their scope remains limited. While methods like [39, 25, 4] have delved into motion planning, visual affordance learning has also gained attention [35, 54, 52, 62, 15]. Other works [7, 58] have proposed special representations for manipulation, though they are tailored mainly for suction grippers. On the other hand, [13, 14] introduce generalizable parts as a novel representation but are limited to single-part interaction and can not handle natural instructions.
Generalizable object manipulation
@@
Achieving generalization in robot applications is essential yet challenging. Various works [8, 47, 18, 12, 57] employ supervised learning with motion planning to tackle generalizable tasks such as grasping. Nevertheless, these techniques often necessitate task-specific architectures and may falter in intricate manipulations. Though reinforcement learning holds promise for complex challenges [40, 2], the lack of generalizability remains unresolved [30, 16]. The ManiSkill benchmark [36] pioneers category-level object manipulation in simulations, while [42] utilizes imitation learning with RL-trained demonstrations. However, in both cases, the generalization ability is still unclear. [13, 14] first tackles generalizable manipulation in a cross-category manner, but is still limited to predefined part classes. [56, 51] have universal generalization ability but only for grasping tasks. In summary, Generalizable object manipulation is still extremely challenging.
Robot control with large language models
@@
Incorporating pre-trained language models into robot systems has been gaining tremendous attention recently. A notable segment of these works concentrates on the grounding of language to manipulation skills, essential for executing long-horizon instructions [26, 46, 34, 5, 41, 1, 17, 19, 43, 21, 59]. A hierarchical approach, seen in another subset of studies, hinges on a two-fold process: establishing a skill set and subsequently strategizing over it using large language models (LLMs) [3, 23, 45, 22, 32, 28]. However, this method is limited by the huge data collection effort, is still constrained to the predefined skill sets, and lacks visual understanding. In contrast, LLMs are employed to spawn code for low-level skills through APIs, though predicting these skills’ continuous parameters remains challenging [31]. Some employ a way-point-based method for efficiency, sidestepping low-level robot actions in favor of trajectory hand poses [44, 17]. But they also require lots of data and have no detailed planning strategy. VoxPoser [24] performs motion planning based on affordance and obstacle maps generated by LLM, but still lacks visual understanding. For articulated objects with diverse geometries, intricate structures, and physical constraints, language-guided manipulation remains a challenging and underexplored problem.
III Problem Formulation
III-A Semantic and Actionable Parts
As shown in Fig 1, we characterize a semantic part based on its human-defined identity or its functional role in articulated objects. During the interaction, for a actionable part, we adopt the concept of ”Generalizable Actionable Part (GAPart)” as formulated by GAPartNet [13]. GAPart categorizes parts across different objects based on their actionability, where parts within the same GAPart category exhibit similar geometry and functionality. For instance, a button designed for pressing and a knob meant for rotation would not fall under the same GAPart category, even if they appear similar. Our approach emphasizes the importance of understanding parts at both semantic and action levels for generalized manipulation of articulated objects.
III-B Problem Definition and Framework Overview
Figure 2: Framework overview. Given a natural language input conditioned on visual observation, we first obtain the semantic parts and interpret the language instructions on them. Then we link the semantic parts to actionable parts through a part grounding module, on which executable actions can be predicted.
Problem definition. Our task focuses on open-vocabulary manipulation of articulated objects. The input consists of a human language instruction and a single-view RGB-D image. A fixed base robot arm is required to follow human instructions and manipulate the articulated object within the scene to achieve the goal specified in the instructions.
Framework overview (Fig. 2). We propose SAGE, a comprehensive framework designed for open-vocabulary articulated object manipulation in the wild. By combining the strength of the general-purpose Visual Language Model (VLM) and a domain-specific 3D part perception and manipulation model, SAGE is uniquely capable of understanding object parts on both semantic and action levels, facilitating more effective and adaptable manipulation in real-world scenarios.
In the SAGE system, we employ the classic perception-decision-execution-feedback loop to generalizable articulate object manipulation: More specifically, (1) we first use GPT-4V to process the input RGB image to obtain a scene description that contains task-related information, such as semantic parts, part interaction possibilities, and object states. We further detect more actionable parts with GAPartNet and fuse both outputs to obtain the final part-aware scene description. (2) Then, we again prompt GPT-4V to comprehend the scene description together with the human instruction to act like a global planner. It outputs action programs step by step and makes decisions based on execution feedback. (3) The action programs are executed with a motion planner which controls the robot arm to follow a predefined trajectory grounded on the actionable part. (4) Finally, we interactively detect the part and object state changes and feedback on the updates to the global planner.
IV Method
IV-A Part-aware Scene Perception
In daily-life scenarios, human language instructions could be diverse and even ambiguous. To facilitate large language models to truly understand human intention and plan accordingly, it is beneficial to add comprehensive scene descriptions to the prompt text inspired by the success of Chain-of-thought [53] prompting.
@@ IV-A Part-aware Scene Perception
Figure 3: Context-aware instruction interpretation. With instruction and observation (RGBD image) as input, the interpreter first generates a scene description with VLM and GAPartNet[13]. Then LLM(GPT-4) takes both instruction and scene description as input and generates semantic parts and action programs. Optionally, we can input a specific user manual, and LLM will generate a target actionable part; see Sec.IV-C for more details about the actionable part.
@@
Figure 4: Scene description. To better help with action generation, the scene description should contain object information, part information, and some interaction-related information. We use expert GAPart model[10] to generate some expert descriptions as part of the VLM prompt and then generate the scene descriptions, which works well and absorbs the advantages of both models.
Scene Description
Chain-of-thought[53] is proven useful for prompting large models. Here, we also notice that, compared to outputting an action program (See Sec. IV-B) directly using a Visual-Language Model, it is better to first generate a scene description $\tilde{D}_{scene}$ containing task-related information (such as parts, object states, interaction possibilities). However, we observe that VLM provides rich context but lacks accuracy in part-related perception as illustrated in Fig. 4. This example illustrates that employing the Visual-Language Model directly can result in the omission of crucial information related to parts and tasks, or may even introduce factual inaccuracies.
Actionable Part Detection
We deploy the 3D GAPart Model [13] to enhance part perception. This model interprets the partial RGBD point cloud $X$ to yield its set of 3D part proposals, which are masked point clouds denoted as $\mathcal{M}^{3D}=\{l_{j},M_{j}^{3D}\}_{j=1}^{n}$, where $l$ is the actionable part label, ${M}$ is the masked partial point cloud, and $j$ indexes the detected actionable parts from $1$ to $n$. We reframe the actionable parts with a template sentence as shown in Fig. 4, eg., There are 4 hinge-door and 4 line-fixed-handle on the object. The sentences that contain actionable parts are then appended to obtain the part-enhanced scene description ${D}_{scene}$. We empirically found this significantly boosts the performance of perception. More details are given in the experiments.
IV-B Instruction Interpretation and Global Planner
Given the part-aware scene description ${D}_{scene}$, we here again employ GPT-4V [37] as the instruction interpreter to translate natural-language instruction $\tilde{I}$ into executable actions that will be sent to the downstream execution module. We devise a representation named action unit to encapsulate the output actions.
Action Units
@@ IV-B Instruction Interpretation and Global Planner
We define the most basic manipulation on one single part of articulated objects as an “action unit”, which can be represented by a 3-tuple ${\bm{a}}=(\tilde{p},j,\Delta s)$ with part name $\tilde{p}$, joint type $j$, and the change of part state $\Delta s$. The part name $\tilde{p}$ is a noun in natural language. $j$ refers to the joint directly linked to part $\tilde{p}$ and indicates whether it is a revolute or a prismatic joint; $\Delta s$ is the state change of joint $j$ under the action, which is either an angle ±$\pm\theta$ if $j$ is revolute or a relative translation ±$\pm t$ w.r.t. the part bounding box if $j$ is prismatic. The final output of the interpreter should be in one of the following formats:
@@
Here the Union expression is for non-deterministic policy generation when multiple action units can reach the same goal, e.g., both pulling the door and pressing a button result in the microwave door being opened. List is for sequential action generation when no single-step solution exists, e.g., to open a door, a knob must be first rotated to unlock the door.
Global Planner
Throughout the interaction, tracking of both the target gripper and part states is maintained by a global planner, where details are given in Algorithm 1. If significant discrepancies arise, the planner has the discretion to select from one of four states: ”continue”, ”transition to the next step”, ”halt and replan”, or ”success”. For instance, if the gripper is set to rotate 60 degrees along a joint, yet the door only opens to 15 degrees, the LLM planner would opt to ”halt and re-plan”. This interaction tracking model ensures that the LLM remains cognizant of developments during interactions, permitting strategy adjustments and, where necessary, recovery from unforeseen setbacks.
IV-C Part Grounding and Execution
@@
An execution module is implemented to act accordingly to the action unit outputted by the global planner. Like any language-guided manipulation task, we first need to ground the action onto the corresponding physical actionable part.
Actionable part grounding
To this end, We first collect a dataset of actionable part features and part labels $\{{F}^{\text{dino}}_{\text{part}},l_{\text{part}}\}$. This dataset contains the pre-defined interaction policies of each typical GAPart. The policies are a series of pre-defined end-effector trajectories to complete a certain motion for a specific part. It serves as our domain-specific expert which is the key to the success of our generalizable system. In particular, we extract a good image feature map $F^{\text{dino}}$ from DINOv2 [38]. Then, for the target actionable part denoted as $p_{j}$, we use its mask to get the part feature $F^{\text{dino}}_{j}$. We first do max-pooling to obtain the part feature and run KNN algorithm to find the grounded actionable part label denoted as $l_{j}$
@@
Trajectory Generation
Once we have grounded the semantic part to the actionable GAPart, we can generate executable manipulations on this part. We first estimate the part pose $P_{j}$ as defined in GAPartNet [13]. We also compute the joint state (part axis and position) and plausible motion direction based on the joint type (prismatic or revolute). Then we generate the part actions according to these estimations. We first predict an initial gripper pose as the primary action. It is followed by a motion induced from a pre-determined strategy defined in GAPartNet [13] according to the part poses and joint states. For example, to open a door with a revolute joint, the starting position can be on the door’s rim or handle, with the trajectory being a circular arc oriented along the door’s hinge.
IV-D Interactive Feedback
Algorithm 1 Global Planner